Statistics from a county public health nurse is provided below. Assist her in completing her report by determining the rate for each vital statistic based on the following information.

if a rate can not be calculated, note what information is missing. Provide the rate or information needed in column J, coinciding with the information on that rate. For this county, the midyear population is 500,000, the estimated female population at that time was 75,000.

*Note that all rates are multiplied by 1,000 except Maternal Mortality is by a factor of 10,000.

. Fetal deaths: (20+ weeks gestation) 42
Answer #1

Total number of population is 500000

Total number of females - 75000

Fetal deaths - 42

Missing data is number of live births ,if it was available then we could find out the fetal mortality rate

By the given data ,we can find out the crude death rate of fetus

Crude death rate = number of deaths of fetus /total mid year population ×1000

= 42 / 500000×1000

= 42/ 500

= 21/250 = 0.084